Suppose GDP in this country is $680 million. Enter the amount for government purchases.

National Income Account

Value

(Millions of dollars)

Government Purchases (GG)
Taxes minus Transfer Payments (TT) 180
Consumption (CC) 300
Investment (II) 180
National Saving (S) = Y - C - G
= I
Private Saving = Y - C - T
=
Public Saving = T - G
=

Based on your calculations, the government is running a budget   .
Answer #1

Answer:

1]

GDP = C + I + G

680 million = 300 million + 180 million + G

G = 680 million - 480 million

G = $ 200 million

2]

National Saving = Private Saving + Public Saving = GDP - C - G = (680 - 300 - 200) million = 180 million

3]

Private Saving = GDP - T - C = (680 - 180 - 300) million = $ 200 million

4]

Public Saving = T - G = (180 - 200) million = -20 million

5]

Since (T - G) < 0, government is running a budget deficit.